• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 08:42:41 on April 1,2008

    Entry number 229317

    keyword=Shift Summary

    Owl Shift starts
    Changed calibration coefficients to match BCM current reading with MCC
    Harp scan done
    Centered beam on BeO
    1 ua on Carbon See Log book entry for photograph of scalers
    DE Rates varied from 15 to 200 KHz, Higher rates correspond to the lower PMTs
    E rates varied from 2 to 25 KHz, Higher rates correspond to the lower PMTs
    1 ua on BeO Rates roughly doubled compared to above
    1 ua No Target, Rates dropped roughly by a factor of 100

    Turned on Raster
    Put in 6 cm LH2 target no collimstor
    High voltage tripped on E and dE PMTs
    Turn on one PMT Count Rate is 800 KHz.
    Perhaps beam may not be aligned well enough with target center.
    Also 2 mm x 2 mm may be too large

    Put C target back in
    Take quasi elastic data in L HRS. Run 2101
    Look at histogram of data along x. See Quasi elastic peak.
    Degauss Q3 Lower current to 0 and raise to 1400 A. Regulate back to original value.
    Take another quasi elastic run on C. Run 2102
    Compare runs 2101 and 2102. They look pretty much identical. Need
    narrower peak for better comparison. Do have time to use sieve slit.

    Reduced beam size to 1 mm x 1 mm. Check on BeO target. Reposition beam Put in C target. Turn on all PMTs and turn on high voltage on wire chambers. See picture. Prescale T1 and T2 to about 40 Hz. Take Run 2104.

    Lower beam current to 0.5 ua. Put in LH2 6 cm target Run 2108
    Deadtime is 50 % even after prescaling T1 abd T2 down to 40 Hz.

    Owl shift ends